How much grams of H3PO3 and Ca(OH)2 are needed for 100 grams of Ca(PO4)2
1
Expert's answer
2020-05-21T13:56:39-0400

Balanced equation of H3PO3 and Ca(OH)2 is

2H3PO3 + 3Ca(OH)2 -----> Ca3(PO3)2 +6H2O

Molar mass of H3PO3 is 82gm.

Molar mass of Ca(OH)2 is 74gm.

Molar mass of Ca3(PO3)2 is 278gm.

Now 2*82=164gm of H3PO3 produces 278gm of Ca3(PO3)2 ,

So, 278gm of Ca3(PO3)2 produced by 164gm of H3PO3

Hence 100gm of Ca3(PO3)2 is produced by 164*100/278= 59gm of H3PO3.

similarly 278gm of Ca3(PO3)2 is produced by 3*74gm of Ca(OH)2

so 100gm of Ca3(PO3)2 is produced by 3*74*100/278 =79.85 = 80gm of Ca(OH)2

Hence 80gm of Ca(OH)2 is produced.